OpenStudy (anonymous):

help please? find d^2y/dx^2 for y+x^4/600, at x=e

OpenStudy (anonymous):

a) 0.02e b)0.02e2 c)0.01e2 d)0.02 e)0.01e

OpenStudy (tkhunny):

Sorry, this one doesn't quite make sense. Do you mean \(y = \dfrac{x^{4}}{600}\)

OpenStudy (anonymous):

yes @tkhunny

OpenStudy (tkhunny):

Okay, then \(\dfrac{dy}{dx} = \dfrac{4x^{3}}{600} = \dfrac{x^{3}}{150}\) You do the next one.